(1.00-.04)(1+r -.0050^6=1.4185
.96(1+r -.005)^6=1.4185
(1+r -.005)^6=1.4776
I know from class the answer is .0722. However, I can not work the problem beyond this point to solve. I believe I need to divide the exponent by 1/6
on both sides of the equation. But when I do, I do not get the correct answer.
My calculator does not have a ^. Even if it does I want to solve the problem by long hand first.

Answer by checkley77(12844)

(1.00-.04)(1+r -.0050^6=1.4185
.96(1+r -.005)^6=1.4185
(1+r -.005)^6=1.4776
r+.995
r+.995
-----------
r^2+.995r+.995r+.990025
r^2+1.99r+.990025
r+.995
------------------------
r^3+1.99r^2+.990025r+.995r^2+1.97r+.985
r^3+2.985r^2+2.96r+.985
r+.995
-----------------------------
[I don't really think you want to go down this road, it's more work than I believe you can handle.
---------------------------------
The following is the straight forward solution using roots.
(r+.995)=6throot(1.4776)
r+.995=1.06723
r=1.06723-.995
r=.0722 ans.